OpenStudy (anonymous):

log[3](2x+5) = log[3]9+2

Nnesha (nnesha):

\[\huge\rm log_3 (2x +5) = \log_3 (9x+2)\] i guess that suppose to be 9x right ??

Nnesha (nnesha):

right side

Nnesha (nnesha):

??

OpenStudy (anonymous):

um no its not

Nnesha (nnesha):

so 9+2 ??

OpenStudy (anonymous):

yes

Nnesha (nnesha):

are u sure ?

OpenStudy (anonymous):

thats what the paper says

Nnesha (nnesha):

\[\huge\rm log_3 (2x +5) = \log_3 (9+2)\] okay then both side have log_3 base so you can cancel out

OpenStudy (anonymous):

ok \[2x+5 = 9+2\]

Nnesha (nnesha):

yes solve for x

OpenStudy (anonymous):

why do the logs cancel out ? my teacher hasn't told us anything abut that i dont believe

Nnesha (nnesha):

bec both side log r same \[\log_3 = \log_3\] what you would get when you divide both side by log_3 ?

OpenStudy (anonymous):

1

Nnesha (nnesha):

yes so 1 (2x+5) = 1(9+2)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

what times are you on here because you are a HUGE help !!

Nnesha (nnesha):

usually 3-9 pm

OpenStudy (anonymous):

ok you will most likely find me on here tomorrow . and is that pacific time?

Nnesha (nnesha):

nope